Word problems inolving ratios and poportions, percen, etc..
50 is what pecent of 80?
x * 80/100 = 50
=> x = 500/8 = 62.5
82 of what number is 16?
x/100 * 82 = 16
=> x = 1600/82=800/41 = 19.51

Answer by sowmya(32)

50 is what pecent of 80?
x * 80/100 = 50
=> x = 500/8 = 62.5
82 of what number is 16?
x/100 * 82 = 16
=> x = 1600/82=800/41 = 19.51